OpenStudy (anonymous):

given that f(100) = 150, f'(100) = -3, find f'(105.5)

OpenStudy (anonymous):

are you sure it isn't f(105.5)? this is like a linear approximation problem

OpenStudy (anonymous):

I was wondering the same thing. You want to get the new f from the adding the slope times the small change in the independent variable to the original f. f(b) = f(a) + (df/dx)(b-a) usually need (b-a) small.

OpenStudy (anonymous):

the formula above is the linear approximation formula, so i'm guessing it is indeed approximating f(105.5) L(105.5) = f(100) + f'(100) (105.5 - 100) = 150 - 3(5.5) = 133.5